OpenStudy (anonymous):

graph the equation and identify the y-concept x+5y=15

OpenStudy (anonymous):

put into y= mx + b form, where m is slope and b is y-intercept: y = -1/5x + 3 So it's a line with slope -1/5 that crosses the y axis at (0,3).

OpenStudy (anonymous):

And it's "y-intercept," not "y-concept."

OpenStudy (anonymous):

y=-x+15/5